How much money should a company deposit in an account with an annual interest rate of 4%, compounded quarterly, in ord...

  1. How much money should a company deposit in an account with an annual interest rate of
    4%, compounded quarterly, in order to have $50,000 in six years? (Round to the nearest dollar)

    1. (a) $41,923

    2. (b) $38,844

    3. (c) $39,378

    4. (d) $50,000
Answer #1
Future Value = C*[(1+(r/m))^mt]
where C is the present value
r is the interest rate that is 4%.
t is the time period in years that is 6.
m is the compounding period that is 4.
Future value = 50000
C*[(1+.04/4))^4*6] = 50000
C*[(1.01)^24] = 50000
C*[1.269735] = 50000
C = 50000/(1.269735)
C = 39378.31.
The company should deposit
(c) $39,378
in the account.